Basal energy requirement. The basal energy requirement B is the number of calories that a person needs to maintain the life process. For a 28-yeaar-old female with a height of 160 centimeters and a weight of 45 kilograms(kg), B is 1300 calories. If her weight increases to 50 kg, then B is 1365 calories. There is a linear equation that expresses B in terms of her weight w. Find the equation and fiind the basal energy requirement if her weight is 53.2 kg.

Answer by stanbon(75887)

Basal energy requirement. The basal energy requirement B is the number of calories that a person needs to maintain the life process. For a 28-yeaar-old female with a height of 160 centimeters and a weight of 45 kilograms(kg), B is 1300 calories. If her weight increases to 50 kg, then B is 1365 calories. There is a linear equation that expresses B in terms of her weight w. Find the equation and fiind the basal energy requirement if her weight is 53.2 kg.
------------------
You have two points that relate weight to B:
(45,1300) and (50,1365)
------------------------
slope = (1365-1300)/(50-45) = 65/5 = 13
Find the intercept:
1300=13*45 + b
b = 715
-------------
EQUATION:
B(x) = 13x + 715
-----------------
Find the equation and fiind the basal energy requirement if her weight is 53.2 kg.
B(53.2)=13*53.2+715
B(54.2) = 1406.60
======================================
Cheers,
Stan H.